### v1

# Close-Talk, and Pseudo-Label Based Far-Field, Speech Enhancement

Abstract
The current dominant approach for neural speech enhancement is via purely-supervised deep learning on simulated pairs of far-field noisy-reverberant speech (i.e., mixtures) and clean speech. The trained models, however, often exhibit limited generalizability to real-recorded mixtures. To deal with this, this paper investigates training enhancement models directly on real mixtures. However, a major difficulty challenging this approach is that, since the clean speech of real mixtures is unavailable, there lacks a good supervision for real mixtures. In this context, assuming that a training set consisting of real-recorded pairs of close-talk and far-field mixtures is available, we propose to address this difficulty via close-talk speech enhancement, where an enhancement model is first trained on simulated mixtures to enhance real-recorded close-talk mixtures and the estimated close-talk speech can then be utilized as a supervision (i.e., pseudo-label) for training far-field speech enhancement models directly on the paired real-recorded far-field mixtures. We name the proposed system ctPuLSE. Evaluation results on the CHiME-4 dataset show that ctPuLSE can derive high-quality pseudo-labels and yield far-field speech enhancement models with strong generalizability to real data.
Index Terms:
Close-talk speech enhancement, pseudo-label based speech enhancement, robust automatic speech recognition.
I Introduction
Dramatic progress has been made in speech enhancement [1], thanks to the rapid development of deep learning. The current dominant approach is based on supervised deep learning, where paired noisy-reverberant speech (i.e., mixtures) and clean speech are simulated and utilized to train deep neural networks (DNN) to predict the clean speech based on its paired mixtures in a purely-supervised, discriminative way [1, 2]. The trained models, however, often exhibit limited generalizability to real-recorded mixtures [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], mainly because the simulated training data is typically, and in many cases inevitably, mismatched with real test data.
To improve the generalizability, this paper investigates training models directly on real-recorded, target-domain mixtures. This approach, however, cannot be straightforwardly realized, since the underlying clean speech is not available for real mixtures, unlike simulated mixtures, where the paired clean speech can be readily available through room simulation and can serve as a fine-grained supervision at the sample level for model training. The key to enable successful training on real mixtures, we think, is to figure out a mechanism that can reliably compute a high-quality supervision signal (i.e., pseudo-label or pseudo-target) for real-recorded mixtures.
One possible way, we propose, is to leverage close-talk mixtures. During data collection, besides using far-field microphones to record target speech, a close-talk microphone is often placed near the target speaker to record the close-talk speech at the same time. See Fig. 1 for an illustration. The recorded close-talk mixture, while the target speaker is talking, usually has a much higher SNR of the target speaker than any far-field mixture, simply due to the very short distance from the target speaker to its close-talk microphone. Although the close-talk mixture is typically not perfectly clean, as non-target signals (such as environmental noises, room reverberation and competing speech) could also be picked up by the close-talk microphone, it usually exhibits a very high input SNR and hence could be utilized to compute a reliable, high-quality supervision for real-recorded far-field mixtures.
Figure 1: Scenario illustration. Close-talk mixture consists of close-talk speech and noises, and far-field mixture consists of far-field speech and noises. Best viewed in color.
With this understanding, this paper first investigates close-talk speech enhancement, a task which aims at enhancing close-talk mixtures and estimating close-talk speech. Solving this task could enable many applications. One of them, which is investigated in this paper, is that, based on the estimated close-talk speech, a pseudo-label can be derived for each real-recorded far-field mixture and used as a supervision to train supervised models directly on real-recorded far-field mixtures, thereby potentially realizing better generalizability to real data. We summarize the contributions of this paper as follows:
  • For the purpose of achieving better far-field speech enhancement, we propose to first investigate close-talk speech enhancement (CTSE), which aims at enhancing close-talk mixtures to estimate close-talk speech. Although CTSE is a particular form of speech enhancement (i.e., in close-talk conditions) and has been studied for other purposes [14, 15, 16], we point out that it is a valuable task that could yield high-quality pseudo-labels for real-recorded data, and is hence worth investigating.
  • We propose ctPuLSE, a pseudo-label based approach for far-field speech enhancement, where estimated close-talk speech is utilized to derive a supervision for training supervised enhancement models directly based on real far-field mixtures.
  • Following SuperM2M [3], an earlier algorithm that trains enhancement models by alternating between supervised and un-/weakly-supervised learning, we propose a co-learning algorithm that trains the same enhancement model using both simulated and real data, where the pseudo-label of real data is derived based on the estimated close-talk speech. This way, the model can also learn from massive amount of simulated training data, which can be easily simulated and can be very helpful when the real training data is scarce. Compared with SuperM2M, we observe that ctPuLSE obtains comparable robust automatic speech recognition (ASR) performance, while much better enhancement performance.
Although ctPuLSE is simple and straightforward, it obtains strong robust ASR and speech enhancement performance on the public CHiME-4 dataset [17], the most popular benchmark to date in robust ASR and speech enhancement. The evaluation results suggest that ctPuLSE can effectively train enhancement models on real-recorded far-field mixtures, and can yield better generalizability to real data than purely-supervised models trained only on simulated data. A sound demo is provided in the link below [^1] See https://zqwang7.github.io/demos/ctPuLSE_demo/index.html.
II Related Work
ctPuLSE is related to earlier works in four key aspects.
II-A Generalizability of Supervised Enhancement Models
Improving the generalizability of supervised learning based speech enhancement models (trained on simulated data) to real-recorded data has received decade-long research interests. The current dominant approach [18, 19, 1, 12, 13] is to simulate massive amount of training data to cover as many variations (that could happen in real-recorded test data) as possible. However, the generalizability is often limited by the current simulation techniques, which usually cannot simulate mixtures as realistic as real-recorded mixtures. This can be observed from recent speech enhancement/separation and robust ASR challenges such as CHiME-{4,5,6,7} [17, 20, 10], AliMeeting [21], MISP [22], and Clarity [8], where using the immediate outputs from DNN-based enhancement or separation models trained on simulated mixtures for robust ASR and human hearing has had limited successes [11, 23]. Different from this stream of research, this paper investigates training enhancement models directly on real-recorded data to improve the generalizability.
II-B Leveraging Close-Talk Mixtures for Speech Enhancement
Leveraging real-recorded close-talk mixtures as a weak supervision to train speech enhancement and speaker separation models directly on real-recorded far-field mixtures has attracted research interests recently. A representative algorithm in this direction is SuperM2M [3], which builds upon a weakly-supervised speaker separation algorithm named mixture to mixture (M2M) [24] and an unsupervised speaker separation algorithm named UNSSOR [25]. M2M trains enhancement models to produce a speech estimate and a noise estimate such that the two estimates can be linearly filtered to recover observed mixtures, and SuperM2M, building upon M2M, leverages supervised learning on simulated mixtures to improve M2M. Although SuperM2M has shown strong potential for neural speech enhancement and robust ASR, we often observe that SuperM2M cannot sufficiently suppress non-target signals, likely because its loss function is defined on reconstructed mixtures (i.e., the summation of linearly-filtered source estimates), rather than directly on source estimates. In comparison, the loss function in ctPuLSE, which we will show, is defined based on pseudo-labels derived from close-talk mixtures. As long as the pseudo-labels are high-quality, we can reasonably expect that the suppression of non-target signals would be more sufficient and aggressive.
On the other hand, there are studies [22] using oracle speaker-activity timestamps and pre-trained models (e.g., DNSMOS [26]) to select segments of close-talk mixtures that are almost clean, and then using the selected segments of mixtures to synthesize far-field mixtures for training supervised learning based models. However, the training data is still simulated, and the models are not trained on real data.
II-C Pseudo-Label Based Speech Enhancement
There have been studies adapting pre-trained speech enhancement models to target domains via pseudo-labeling. RemixIT [6], a representative algorithm in this direction, first uses a pre-trained enhancement model (named teacher) to enhance target-domain real-recorded mixtures (and generate pseudo-labels), and then another enhancement model (named student) is trained in a supervised way to estimate the generated pseudo-labels. The teacher and student models are designed to update continuously and iteratively to gradually improve each other. Another recent study in this direction is SSST [27], which, building upon RemixIT, introduces an adversarial training algorithm to learn domain-invariant hidden representations and proposes a data-selection mechanism to pick a subset of target-domain mixtures whose pseudo-labels are sufficiently reliable for RemixIT-style training.
Similarly to those studies, ctPuLSE leverages pseudo-labeling as well, but it has a very different problem setup, where paired close-talk and far-field mixtures are assumed available for model training. In this setup, much better pseudo-labels could potentially be computed for far-field mixtures, simply due to the innate high input SNR of close-talk mixtures. This potential could make ctPuLSE a more attractive solution for practical product development, as long as the paired close-talk mixtures can be recorded while collecting far-field mixtures for training.
II-D Close-Talk Speech Enhancement
Our study leverages close-talk speech enhancement to derive high-quality pseudo-labels for far-field speech enhancement. There are existing studies [14, 15, 16] on close-talk speech enhancement, but for different purposes. For example, in [14, 15], the task is to enhance the target speech captured by a close-talk microphone with the help of an ear-mounted microphone; and in [16], the task is to enhance close-talk speech when users talk to dual-microphone mobile phones at a very short distance. Their purpose is not on training models on real data and improving the generalizability to real data.
III Physical Model and Approach Overview
In a noisy-reverberant environment with a $P$-microphone far-field microphone array and a single target speaker wearing a close-talk microphone (see Fig. 1 for an illustration), the physical model of the recorded close-talk mixture and each far-field mixture can be respectively formulated, in the short-time Fourier transform (STFT) domain, as follows:
$Y_{0}(t,f)=X_{0}(t,f)+V_{0}(t,f),$ (1)
$Y_{p}(t,f)=X_{p}(t,f)+V_{p}(t,f),$ (2)
where the subscript $p\in\{1,\dots,P\}$ indexes the $P$ far-field microphones, subscript 00 indexes the close-talk microphone, $t\in\{0,\dots,T-1\}$ indexes $T$ frames, and $f\in\{0,\dots,F-1\}$ indexes $F$ frequency bins. In (1), $Y_{0}(t,f)$, $X_{0}(t,f)$, and $V_{0}(t,f)$ respectively denote the STFT coefficients of the close-talk mixture, close-talk speech, and non-speech signals (such as environmental noises) captured by the close-talk microphone at time $t$ and frequency $f$. Similarly, in (2), $Y_{p}(t,f)$, $X_{p}(t,f)$, and $V_{p}(t,f)$ are respectively the STFT coefficients of the far-field mixture, far-field speech, and non-speech signals captured by far-field microphone $p$ at time $t$ and frequency $f$. In the rest of this paper, when dropping the indices $p$, $t$ and $f$, we refer to the corresponding spectrograms. In this paper, $V$ is assumed to contain an unknown number of strong, non-stationary diffuse and directional noises sources.
Figure 2: ctPuLSE illustration. A CTSEnet is first trained on simulated mixtures for close-talk speech enhancement. It is then used to enhance real-recorded close-talk mixtures and obtain estimated close-talk speech. Next, a ctPuLSEnet is trained for far-field speech enhancement, based on the estimated close-talk speech and its paired real-recorded far-field mixtures.
Fig. 2 illustrates the proposed ctPuLSE algorithm. Given a set of simulated pairs of clean speech and mixtures as well as a set of real-recorded pairs of close-talk and far-field mixtures for training, the proposed system consists of three steps: (a) we train a speech enhancement model (denoted as CTSEnet) based on the simulated training mixtures for close-talk speech enhancement; (b) we apply the trained CTSEnet to enhance each real-recorded close-talk mixture and obtain an estimated close-talk speech; and (c) we leverage the estimated close-talk speech as pseudo-labels to train a far-field speech enhancement model (denoted as ctPuLSEnet) on the paired real-recorded far-field mixtures.
Next, we detail CTSEnet and ctPuLSEnet.
IV CTSEnet
Based on a set of simulated training mixtures, we train CTSEnet via single-microphone complex spectral mapping [28, 29] to predict the real and imaginary (RI) components of target speech $X_{a}$ based on the RI components of the input mixture $Y_{a}$, where, depending on the application scenarios, the microphone index $a$ can index all or a subset of the simulated microphones in the training data. Following [30], we define the loss function on the RI components and magnitude of the DNN estimate $\hat{X}_{a}$:
$\mathcal{L}_{X,a}^{\text{simu}}=\frac{\sum\nolimits_{t,f}\mathcal{F}\Big{(}\hat{X}_{a}(t,f),X_{a}(t,f)\Big{)}}{\sum\nolimits_{t^{\prime},f^{\prime}}\big{|}X_{a}(t^{\prime},f^{\prime})\big{|}},$ (3)
$\mathcal{F}\Big{(}\hat{X}_{a}(t,f),X_{a}(t,f)\Big{)}=$ $\Big{|}\mathcal{R}\big{(}\hat{X}_{a}(t,f)\big{)}-\mathcal{R}\big{(}X_{a}(t,f)\big{)}\Big{|},$
$+\Big{|}\mathcal{I}\big{(}\hat{X}_{a}(t,f)\big{)}-\mathcal{I}\big{(}X_{a}(t,f)\big{)}\Big{|},$
$+\Big{|}|\hat{X}_{a}(t,f)|-|X_{a}(t,f)|\Big{|},$ (4)
where ||$\cdot$|| ⋅ | computes the magnitude or absolute value, $\mathcal{R}(\cdot)$ and $\mathcal{I}(\cdot)$ respectively extract the RI components, and the denominator in (3) balances the loss values across different training mixtures. Other configurations of CTSEnet are detailed later in Section VI-C and VI-D.
V ctPuLSEnet
Once a CTSEnet has been trained, we apply it to enhance each monaural real-recorded close-talk mixture to obtain an estimate of the close-talk speech, denoted as $\hat{X}_{0}^{\text{CTSE}}$, which is then leveraged as a pseudo-label to train ctPuLSEnet on real-recorded far-field mixtures. This section describes the DNN configurations, loss functions, and a co-learning algorithm that trains ctPuLSEnet on both simulated and real mixtures.
V-A DNN Configurations
ctPuLSEnet is trained via single- or multi-microphone complex spectral mapping [31, 32, 33] to predict the RI components of target speech at a designated reference microphone $q\in\{1,\dots,P\}$ based on the RI components of stacked input mixtures. In monaural cases, the DNN is trained to predict the target speech $X_{q}$ based on input mixture $Y_{q}$, and in multi-channel cases, it is trained to predict $X_{q}$ based on all the input mixtures stacked in a fixed microphone order. We denote the DNN estimate as $\hat{X}_{q}$. Besides predicting $X_{q}$, the DNN can also be trained to additionally predict non-speech signals $V_{q}$. We observe that this form of multi-task learning can improve enhancement and robust ASR in our experiments.
Next, we describe the loss functions for speech estimate $\hat{X}_{q}$ in Section V-B and for noise estimate $\hat{V}_{q}$ in V-D. Other DNN setups are detailed in Section VI-C and VI-D.
V-B Loss Functions Based on Estimated Close-Talk Speech
Assuming that the close-talk microphone and far-field microphones are approximately time-synchronized, we directly leverage $\hat{X}_{0}^{\text{CTSE}}$ as the pseudo-label for each far-field mixture $Y_{p}$. Since $\hat{X}_{0}^{\text{CTSE}}$ is time- and gain-aligned to close-talk speech $X_{0}$ rather than to far-field speech $X_{p}$ in far-field mixture $Y_{p}$, special care is needed to account for the time delay and gain differences between $\hat{X}_{0}^{\text{CTSE}}$ and $X_{p}$. In this context, we propose to first linearly filter the DNN estimate $\hat{X}_{p}$ at each frequency to align it to $\hat{X}_{0}^{\text{CTSE}}$ before loss computation:
$\mathcal{L}_{X}^{\text{real}}=\frac{\sum\nolimits_{t,f}\mathcal{F}\Big{(}\hat{X}_{p\rightarrow 0}(t,f),\hat{X}_{0}^{\text{CTSE}}(t,f)\Big{)}}{\sum\nolimits_{t^{\prime},f^{\prime}}\big{|}\hat{X}_{0}^{\text{CTSE}}(t^{\prime},f^{\prime})\big{|}},$ (5)
$\hat{X}_{p\rightarrow 0}(t,f)=\mathbf{\hat{g}}_{p}(f)^{\mathsf{H}}\widetilde{\hat{\mathbf{X}}}_{p}(t,f)$ (6)
where $\widetilde{\hat{\mathbf{X}}}_{p}(t,f)=\big{[}\hat{X}_{p}(t-I+1,f),\dots,\hat{X}_{p}(t,f),\dots,\hat{X}_{p}(t+J,f)\big{]}\in{\mathbb{C}}^{I+J}$ stacks a window of T-F units, $\mathbf{\hat{g}}_{p}(f)\in{\mathbb{C}}^{I+J}$ denotes an estimated multi-tap linear filter to be described in (7), and $\mathcal{F}(\cdot,\cdot)$ is a distance metric defined in (IV). Following the forward convolutive prediction (FCP) algorithm [34], we estimate $\mathbf{g}_{p}(f)$ by linearly projecting the DNN estimate $\hat{X}_{p}(\cdot,f)$ to pseudo-label $\hat{X}_{0}^{\text{CTSE}}(\cdot,f)$ at each frequency $f$:
$\hat{\mathbf{g}}_{p}(f)=\underset{\mathbf{g}_{p}(f)}{\text{argmin}}\sum\limits_{t}\Big{|}\hat{X}_{0}^{\text{CTSE}}(t,f)-\mathbf{g}_{p}(f)^{\mathsf{H}}\widetilde{\hat{\mathbf{X}}}_{p}(t,f)\Big{|}^{2}.$ (7)
This is a quadratic problem, which has a closed-form solution. Next, we plug the closed-form solution into (6), compute the loss in (5), and train the DNN.
An alternative way for linear filtering is to estimate the filter in the time domain, following the ideas behind CI-SDR [35]. In detail, we compute a time-domain multi-tap Wiener filter to align $\hat{x}_{p}$ to $\hat{x}_{0}^{\text{CTSE}}$ (where $\hat{x}_{p}=\text{iSTFT}(\hat{X}_{p})$ is the re-synthesized time-domain signal of $\hat{X}_{p}$ obtained via inverse STFT (iSTFT), and similarly $\hat{x}_{0}^{\text{CTSE}}=\text{iSTFT}(\hat{X}_{0}^{\text{CTSE}})$) by solving the following problem:
$\hat{h}_{p}=\underset{h_{p}}{\text{argmin}}\sum_{n}\Big{|}\hat{x}_{0}^{\text{CTSE}}[n]-\big{(}h_{p}*\hat{x}_{p}\big{)}[n]\Big{|}^{2},$ (8)
where $n$ indexes time-domain samples, * denotes linear convolution, and $h_{p}\in{\mathbb{R}}^{K+1+K}$ denotes a $(K+1+K)$-tap linear filter with $K$ past and $K$ future taps. (8) is also a quadratic linear regression problem, where a closed-form solution can be readily computed. After that, we compute the STFT spectrogram of the filtered time-domain DNN estimate
$\hat{X}_{p\rightarrow 0}=\text{STFT}\big{(}\hat{h}_{p}*\hat{x}_{p}\big{)},$ (9)
which is then plugged into (5) to compute the loss for training.
V-C Addressing Time-Synchronization Issues in Pre-Processiing
In the previous subseciton, when performing linear filtering for loss computation, we assume that far-field and close-talk microphones are reasonably time-synchronized. In practical data collection, the close-talk microphone and far-field microphone array are usually managed and processed by two different devices. Although the microphones on each device are typically synchronized, there could be synchronization errors between the microphones on different devices. Later in Section VI-B, we will introduce a technique based on the classic GCC-PHAT algorithm [36, 37] to roughly synchronize the close-talk microphone and far-field array. This technique is utilized as a pre-processing operation prior to training.
V-D Co-Learning on Simulated and Real Mixtures
In practical application scenarios, the amount of real data is typically scarce, as it is often effort-consuming to collect real-recorded close-talk and far-field mixture pairs. In this case, even if ctPuLSEnet can be trained on real data, the performance is often limited, simply due to the limited amount of training data. On the other hand, simulated data can be easily and massively generated through simulation. In this case, following the SuperM2M algorithm [3] which combines supervised learning on simulated data and weakly-supervised learning on real data, we propose to train ctPuLSEnet on both simulated and real data. See Fig. 3 for an illustration, where supervised learning on simulated data is shown in Fig. 3(b).
Figure 3: Illustration of co-learning, which trains the same model via supervised learning based on (a) real-recorded far-field mixtures and their pseudo-labels derived from CTSEnet; and (b) simulated mixtures.
Specifically, at each training step, we randomly sample either a mini-batch of real far-field mixtures and their pseudo-labels computed by CTSEnet or a mini-batch of simulated far-field mixtures (where the clean speech is available) to train the same DNN. If the sampled mini-batch is real, the loss function can be the $\mathcal{L}_{X,q}^{\text{real}}$ loss defined in (5), while if the sampled mini-batch is simulated, the loss can be $\mathcal{L}_{X,q}^{\text{simu}}$ in (3).
Alternatively, we can train ctPuLSEnet to not only predict target speech but also non-target signals via multi-task learning, as is described earlier in Section V-A. In this case, the loss on simulated data can be
$\mathcal{L}_{q}^{\text{simu}}=\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q},$ (10)
where $\mathcal{L}_{V,q}^{\text{simu}}$ denotes the loss on estimated noise $\hat{V}_{q}$ and is defined, by following $\mathcal{L}_{X,q}^{\text{simu}}$ in (3), as
$\mathcal{L}_{V,q}^{\text{simu}}=\frac{\sum\nolimits_{t,f}\mathcal{F}\Big{(}\hat{V}_{q}(t,f),V_{q}(t,f)\Big{)}}{\sum\nolimits_{t^{\prime},f^{\prime}}\big{|}V_{q}(t^{\prime},f^{\prime})\big{|}},$ (11)
and $\mathcal{L}_{Y,q}$ is a mixture-constraint loss [38] encouraging the speech and noise estimates to sum up to the observed mixture:
$\mathcal{L}_{Y,q}=\frac{\sum\nolimits_{t,f}\mathcal{F}\Big{(}\hat{X}_{q}(t,f)+\hat{V}_{q}(t,f),Y_{q}(t,f)\Big{)}}{\sum\nolimits_{t^{\prime},f^{\prime}}\big{|}Y_{q}(t^{\prime},f^{\prime})\big{|}}.$ (12)
Accordingly, the loss on real data can be modified to
$\mathcal{L}_{q}^{\text{real}}=\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}.$ (13)
where $\mathcal{L}_{X,q}^{\text{real}}$ is defined in (5).
Notice that in each of the loss functions in (3), (5), (11) and (12), a normalization term is used in the denominator to balance the loss with the others. This is particularly useful for $\mathcal{L}_{X}^{\text{real}}$ in (5), as the close-talk speech estimated based on the close-talk mixture (i.e., $\hat{X}_{0}^{\text{CTSE}}$) could have a gain level very different from far-field mixtures.
We can also add a weighting term $\alpha\in{\mathbb{R}}_{>0}$ between the loss on simulated data and the loss on real data to balance their importance. In detail, the overall loss is defined as
$\mathcal{L}_{q}=\left\{\begin{aligned} &\alpha\times\mathcal{L}_{q}^{\text{simu}},\text{for mini-batches of simulated data}\\ &\mathcal{L}_{q}^{\text{real}},\text{for mini-batches of real data}\end{aligned}\right.$ (14)
VI Experimental Setup
Figure 4: Pipeline of ASR evaluation for (a) close-talk speech enhancement; and (b) far-field speech enhancement. Enhanced speech (e.g., $\hat{x}_{0}=\text{iSTFT}(\hat{X}_{0})$ in close-talk speech enhancement) is directly fed to a pre-trained backend ASR system for recognition. No joint training between ASR and enhancement models is performed. In (b), an optional speaker reinforcement module [39], which adds a scaled version of input mixture $y_{q}$ to $\hat{s}_{q}$, can be included.
Based on real-recorded mixtures, the main objective of our experiments is to demonstrate whether ctPuLSEnet can achieve better enhancement on real-recorded far-field mixtures than purely-supervised learning based approaches, which can only train enhancement models on simulated data. Bearing this objective in mind, we validate the proposed algorithms on the CHiME-4 dataset [40], which contains both real-recorded and simulated mixtures for both training and testing.222 Besides CHiME-4, we checked many other datasets for evaluation, but CHiME-4 is the only one with both real and simulated mixtures, paired close-talk and far-field mixtures, a setup for ASR evaluation, and a single target speaker in each recording. Closest to CHiME-4 are datasets such as CHiME-{5,6,7} [41, 10], AliMeeting [21] and AMI [42]. However, they are multi-speaker in nature and hence not very suitable for speech enhancement. Since, for real mixtures, the clean speech is not available for evaluation, we mainly check whether the enhanced speech can yield better ASR performance, considering that ASR scores can indicate the degrees of speech distortion and can reflect speech intelligibility. The ASR evaluation pipeline is shown in Fig. 4, where enhanced close-talk or far-field speech is directly fed to a strong pre-trained ASR system for recognition. The rest of this section describes the CHiME-4 dataset, synchronization of close-talk and far-field mixtures, setup for ASR evaluation, miscellaneous system configurations, baselines for comparison, and evaluation metrics.
VI-A CHiME-4 Dataset
CHiME-4 [40, 17, 43] is a major benchmark for evaluating far-field speech recognition and enhancement algorithms. The far-field recording device is a tablet mounted with six microphones, with the second microphone placed on the rear and the other five facing front. During data collection, the target speaker hand-holds the tablet and reads text prompts shown on the screen of the tablet. The target speaker wears a close-talk microphone so that a monaural close-talk mixture can be recorded along with each six-channel far-field mixture.
The mixtures are recorded in four representative daily environments (including buses, cafeteria, pedestrian areas and streets), where multiple strong, non-stationary directional and diffuse noises can naturally exist. In CHiME-4, the room reverberation is weak, and the major challenge is in how to deal with the multi-source non-stationary noise signals.
TABLE I: Number of Utterances in CHiME-4
Close-talk/
Type Far-field? #mics Training Set Validation Set Test Set
SIMU Far-field 6 $7,138$ ($\sim$15.115.115.115.1 h) $1,640$ ($\sim$2.92.92.92.9 h) $1,320$ ($\sim$2.32.32.32.3 h)
SIMU Close-talk - N/A N/A N/A
REAL Far-field 6 $1,600$ ($\sim$2.72.72.72.7 h) $1,640$ ($\sim$2.72.72.72.7 h) $1,320$ ($\sim$2.22.22.22.2 h)
REAL Close-talk 1 $1,600$ ($\sim$2.72.72.72.7 h) $1,640$ ($\sim$2.72.72.72.7 h) $1,320$ ($\sim$2.22.22.22.2 h)
The number of mixtures in CHiME-4 is summarized in Table I. Besides real mixtures, CHiME-4 also provides simulated far-field mixtures for training and testing. We emphasize that, for each real mixture in the training set, in total it has $7$ channels (i.e., $1$ close-talk plus $6$ far-field microphones), while, for each simulated training mixture, it is far-field, $6$-channel, and does not have the paired close-talk mixture.
The sampling rate is $16$ kHz.
VI-B Synchronization of Close-talk and Far-field Mixtures
In CHiME-4, we observe that the far-field mixtures are reasonably synchronized with each other, while significant synchronization errors exist between close-talk and far-field mixtures. For some utterances, the time delay between the close-talk and far-field mixtures can be as large as $0.05$ seconds, which, if correct, means that the distance between the close-talk microphone and far-field array can be as large as $\sim$17171717 meters (i.e., $0.05\times 340$, assuming that the speed of sound in the air is $340$ meters per second). This clearly does not make sense in the CHiME-4 setup, as the speaker hand-holds the tablet while talking. The root cause of the synchronization errors, we think, is that the close-talk microphone and far-field array in CHiME-4 are placed on, and processed by, two different devices.
If the synchronization errors between the close-talk and far-field mixtures are not properly addressed before using them to train ctPuLSEnet, the loss functions in (5) would be less effective, since the hypothesized filter taps of $\mathbf{g}_{p}(f)$ in (7) or $h_{p}$ in (8), which are hyper-parameters shared by all the training utterances, may not be able to compensate the synchronization errors for every training utterance. To deal with this, we could use a very long linear filter to cover the maximum synchronization error of all the training mixtures. This is however problematic, as the longer the filter is, the more likely that the filter can filter any DNN estimate (even if the estimate is a random signal) to approximate the estimated close-talk speech.
To mitigate the synchronization errors, we design an algorithm, motivated by the classic GCC-PHAT algorithm [36, 37], to approximately align the close-talk mixture to far-field mixtures. We emphasize that this is a pre-processing step at the very beginning of our proposed system (i.e., before training).
In detail, for each of the close-talk and far-field mixtures, we consider its magnitude in each frequency as a one-dimensional sequence, and find an integer frame shift $\hat{d}\in{\mathbb{Z}}$ (shared by all the frequencies) that can, at every frequency, best align the magnitude sequence of the close-talk mixture to those of the far-field mixtures. Specifically, let $M_{p}(f)=|Y_{p}(\cdot,f)|\in{\mathbb{R}}^{T}$ denotes the magnitude sequence at frequency $f$ for microphone $p\in\{0,1,\dots,P\}$, and $R_{p}(f)=\text{FFT}\big{(}M_{p}(f)\big{)}\in{\mathbb{C}}^{T}$ denotes the FFT coefficients after applying a $T$-point faster Fourier transform (FFT) to the magnitude sequence. At each frequency $f$, we first compute the GCC-PHAT coefficients between the magnitude sequence of the close-talk microphone and that of a far-field microphone $p$ by
GCC-PHAT (${}_{p}(t,f,d)$
$=\mathcal{R}\Big{(}\frac{R_{0}(t,f)\times R_{p}(t,f)^{*}}{|R_{0}(t,f)|\times|R_{p}(t,f)^{*}|}e^{-j\times 2\pi\frac{t}{T}d}\Big{)}$
$=\text{cos}\Big{(}\angle R_{0}(t,f)-\angle R_{p}(t,f)-2\pi\frac{t}{T}d\Big{)},$ (15)
where $j$ denotes the imaginary unit, $d$ a hypothesized frame delay, and $\mathcal{R}(\cdot)$ extracts the real component. We then enumerate a set of hypothesized frame delays and find a delay that can produce the largest summation of the GCC-PHAT coefficients at all the far-field microphones and T-F units:
$\hat{d}=\underset{d\in\Omega}{\text{argmin}}\sum_{p=1}^{P}\sum_{t=0}^{T-1}\sum_{f=0}^{F-1}\text{GCC-PHAT}_{p}(t,f,d),$ (16)
where $\Omega\in\{-D,\dots,0,\dots,D\}$ is a set of candidate frame delays with $D$, a tunable hyper-parameter, denoting a hypothesized maximum delay on each side. If the resulting best time delay $\hat{d}$ is positive, we advance the close-talk mixture by $\hat{d}$ frames (and pad zeros to the right), and delay it by $\hat{d}$ frames otherwise (and pad zeros to the left).
Notice that this algorithm is designed to compute an integer number of frame shift for each close-talk mixture. In our experiments, for the STFT configuration of (15) and (16), we set the window size to $16$ ms and hop size to $1$ ms (please do not confuse this STFT configuration used in the pre-processing stage with that in the subsequent DNN training stage). A small STFT hop size is used here, as our aim is to roughly align close-talk and far-field mixtures in this pre-processing stage.
We highlight that the alignment is performed at the granularity of $1$ ms. Although sample-level synchronization is definitely desired, it would dramatically increase the amount of computation, as the enumeration would be conducted at the granularity of samples and this computation would become intolerable when the candidate time delay (and time advance) can be as large as $0.05$ second in CHiME-4. In addition, we may not really need accurate sample-level synchronization, since the STFT window size of our enhancement models can be as large as $32$ ms and the linear filtering in, e.g., (6) could account for slight synchronization errors (that are sufficiently smaller than the $32$ ms window size).
VI-C Training Setup
For CTSEnet, it is trained based on all the $7,138\times 6$ monaural simulated mixtures. Following [3], we apply an SNR augmentation technique to the simulated training mixtures of CHiME-4. That is, during training, we optionally modify the SNR of each simulated mixture, on the fly, by $u$ dB, with $u$ uniformly sampled from the range $[-10,+15]$ dB. No other data augmentation is used.
For ctPuLSEnet, we use all the $7,138$ simulated and $1,600$ real utterances for training. For monaural ctPuLSEnet, we train it on all the $(7,138+1,600)\times 6$ monaural mixtures. For $2$-channel ctPuLSEnet, at each training step we sample two microphones from the front five microphones as input, and ctPuLSEnet is trained to predict the target speech at the first of the two selected microphones. For $6$-channel ctPuLSEnet, it stacks all the six microphones in a fixed order as input to predict the target speech at the fifth microphone. We apply the same SNR augmentation used in CTSEnet to the simulated training mixtures when training ctPuLSEnet. We always train ctPuLSEnet on the combination of the simulated and real mixtures using the co-learning algorithm introduced in Section V-D, as there are only $1,600$ real mixtures (which amount to only $\sim$2.72.72.72.7 hours) in the training set of CHiME-4
For simplicity, we did not filter out microphone signals with any microphone failures in training and inference. We expect ctPuLSEnet to learn to robustly deal with the failures, as it is trained on real mixtures.
VI-D Miscellaneous System Configurations
For STFT, the window and hop sizes are respectively $32$ and $8$ ms, and the square root of Hann window is used as the analysis window.
TF-GridNet [38] is employed as the DNN architecture. It has shown strong separation and enhancement performance recently in serveral representitive supervised speech separation benchmarks. Following the symbols defined in Table I of [38], we set its hyper-parameters to $D=128$, $B=4$, $I=1$, $J=1$, $H=200$, $L=4$ and $E=4$ (please do not confuse the symbols with the ones defined in this paper). The model has $\sim$5.45.45.45.4 million trainable parameters. This configuration of TF-GridNet is used for both CTSEnet and ctPuLSEnet.
We train both CTSEnet and ctPuLSEnet on eight-second segments using a mini-batch size of one. Adam is used as the optimizer. The learning rate starts from $0.001$ and is halved if the loss is not improved in two epochs.
VI-E ASR Model
For both close-talk and far-field speech enhancement, we check whether they can result in better ASR performance by directly feeding their enhanced speech to a pre-trained ASR model for decoding, following the evaluation pipeline shown in Fig. 4(a) and (b).
The ASR model is pre-trained in a multi-conditional way on the official CHiME-4 simulated and real mixtures plus the clean speech signals in WSJ00 by using the script proposed in [44] [^3] https://github.com/espnet/espnet/blob/master/egs2/chime4/asr1/conf/tuning/train_asr_transformer_wavlm_lr1e-3_specaug_accum1_preenc128_warmup20k.yaml, which is available in the ESPnet toolkit. It is an encoder-decoder- and transformer-based system, trained on WavLM features [45] and using a transformer-based language model for decoding. From the results reported in [44] and [46], this model is the current best ASR model on CHiME-4.
We have successfully reproduced the ASR system in [44]. The mixture ASR results (shown later in row 00 of Table III) are very close to the ones reported in row $7$ of Table $1$ of [44].
VI-F Evaluation Metrics
For real-recorded mixtures, where the corresponding clean speech is unavailable for evaluation, we use word error rates (WER) as the major evaluation metric. WER can partially reflect the intelligibility of enhanced speech. Besides WER, DNSMOS [26] is employed to evaluate the quality of enhanced speech.
For simulated mixtures, where the clean speech is available, our evaluation metrics include short-time objective intelligibility (STOI) [47], wide-band perceptual evaluation of speech quality (WB-PESQ) [48], signal-to-distortion ratio (SDR) [49], and scale-invariant SDR (SI-SDR) [50]. They are designed to evaluate the intelligibility, quality, and accuracy of the magnitude and phase of enhanced speech. They are widely-adopted metrics in speech enhancement.
VI-G Baselines
A major baseline is purely-supervised speech enhancement models trained only on the CHiME-4 simulated data. That is, we only use Fig. 3(b) for model training, and use exactly the same training setup as that in ctPuLSEnet. On the other hand, since CHiME-4 is a popular public dataset, we can compare our ASR and enhancement results with the ones obtained by many earlier studies.
VI-H Tricks to Improve ASR Performance
At run time, in default we feed $\hat{x}_{q}$ produced by ctPuLSEnet for ASR decoding. Alternatively, following [3] we apply an existing technique named speaker reinforcement [39] to mitigate speech distortion incurred by enhancement (see Fig. 4(b) for an illustration). It adds a scaled version of the mixture signal $y_{q}$ back to enhanced speech $\hat{x}_{q}$ before performing ASR decoding. That is, the signal sent for ASR is $\hat{x}_{q}+\eta\times y_{q}$, where $\eta$ is computed such that $10\times\text{log}_{10}\big{(}\|\hat{x}_{q}\|_{2}^{2}/\|\eta\times y_{q}\|_{2}^{2}\big{)}=\gamma$ dB. This technique has been found effective at improving ASR performance in [39, 3].
VII Evaluation Results
VII-A Results of Close-Talk Speech Enhancement
Table II presents the results of monaural CTSEnet.
Row 00 reports the results of unprocessed close-talk mixtures. We observe that the ASR results are very good (e.g., $1.49\%$ on the real test set), indicating that the close-talk mixtures already have very high input SNRs and that microphone failures in the close-talk mixtures are minimal. In comparison, the DNSMOS scores are not good (e.g., $2.50$ on the real test set), indicating that non-speech signals captured along with close-talk speech are still significant. These non-speech signals need to be removed in order to derive high-quality pseudo-labels for far-field mixtures.
TABLE II: Results on CHiME-4 Close-Talk Mixtures
[…]
Training data Val. Test Val. Test
Row Systems REAL REAL REAL REAL
0 - Close-talk mixture 2.818 349 123 916 3092.8183491239163092.818\,349\,123\,916\,3092.818 349 123 916 309 2.502 779 209 488 744 42.50277920948874442.502\,779\,209\,488\,744\,42.502 779 209 488 744 4 1.143 109 9961.1431099961.143\,109\,9961.143 109 996 1.490 027 5581.4900275581.490\,027\,5581.490 027 558
1 S CTSEnet 3.263 852 279 278 555 43.26385227927855543.263\,852\,279\,278\,555\,43.263 852 279 278 555 4 3.216 589 413 447 3233.2165894134473233.216\,589\,413\,447\,3233.216 589 413 447 323 1.150 484 899 885 6891.1504848998856891.150\,484\,899\,885\,6891.150 484 899 885 689 1.555 420 617 497 314 31.55542061749731431.555\,420\,617\,497\,314\,31.555 420 617 497 314 3
Row $1$ reports the results of CTSEnet trained via supervised learning on the CHiME-4 simulated mixtures (denoted as S in the “Training data” column). Compared with row 00, the DNSMOS scores are clearly improved (e.g., from $2.50$ to $3.22$ on the real test set). The ASR performance becomes worse, possibly because CTSEnet is trained only on simulated data and could hence introduce some distortion to target speech. Nonetheless, the ASR performance degrades rather slightly (e.g., from $1.49$% to $1.56$% WER on the real test set), meaning that the estimated close-talk speech is still of high quality and could be a good pseudo-label for far-field mixtures.
With a grain of salt, the ASR results in row $1$ can be viewed as the upper-bound performance of ctPuLSEnet.
TABLE III: ctPuLSE vs. Purely-Supervised Models on CHiME-4 Far-Field Mixtures (Single-Channel Input)
[…]
0 Mixture - - - - - 7.506 217 692 752 9057.5062176927529057.506\,217\,692\,752\,9057.506 217 692 752 905 7.539 180 193 284 8817.5391801932848817.539\,180\,193\,284\,8817.539 180 193 284 881 1.273 397 988 803 4591.2733979888034591.273\,397\,988\,803\,4591.273 397 988 803 459 0.870 243 137 103 452 40.87024313710345240.870\,243\,137\,103\,452\,40.870 243 137 103 452 4 1.520 743 794 895 107 71.52074379489510771.520\,743\,794\,895\,107\,71.520 743 794 895 107 7 1.392 663 582 045 257 81.39266358204525781.392\,663\,582\,045\,257\,81.392 663 582 045 257 8 5.932 890 8555.9328908555.932\,890\,8555.932 890 855 4.070 946 5684.0709465684.070\,946\,5684.070 946 568 8.288 195 7418.2881957418.288\,195\,7418.288 195 741 4.470 082 6754.4700826754.470\,082\,6754.470 082 675
1a Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ 17.295 841 213 880 16317.29584121388016317.295\,841\,213\,880\,16317.295 841 213 880 163 17.665 333 817 529 917.665333817529917.665\,333\,817\,529\,917.665 333 817 529 9 2.361 246 844 584 2922.3612468445842922.361\,246\,844\,584\,2922.361 246 844 584 292 0.959 758 057 782 008 20.95975805778200820.959\,758\,057\,782\,008\,20.959 758 057 782 008 2 3.327 439 949 427 8173.3274399494278173.327\,439\,949\,427\,8173.327 439 949 427 817 3.207 257 983 124 223.207257983124223.207\,257\,983\,124\,223.207 257 983 124 22 3.488 200 589 970 5013.4882005899705013.488\,200\,589\,970\,5013.488 200 589 970 501 2.182 971 348 501 050 82.18297134850105082.182\,971\,348\,501\,050\,82.182 971 348 501 050 8 7.634 478 894 284 6467.6344788942846467.634\,478\,894\,284\,6467.634 478 894 284 646 5.236 115 652 295 763 55.23611565229576355.236\,115\,652\,295\,763\,55.236 115 652 295 763 5
1b Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}$ 17.229 124 567 725 44217.22912456772544217.229\,124\,567\,725\,44217.229 124 567 725 442 17.647 178 119 958 08817.64717811995808817.647\,178\,119\,958\,08817.647 178 119 958 088 2.489 149 991 310 0042.4891499913100042.489\,149\,991\,310\,0042.489 149 991 310 004 0.962 018 627 202 278 70.96201862720227870.962\,018\,627\,202\,278\,70.962 018 627 202 278 7 3.317 335 051 347 744 63.31733505134774463.317\,335\,051\,347\,744\,63.317 335 051 347 744 6 3.211 204 003 775 436 43.21120400377543643.211\,204\,003\,775\,436\,43.211 204 003 775 436 4 3.392 330 383 480 8263.3923303834808263.392\,330\,383\,480\,8263.392 330 383 480 826 2.087 097 606 843 912.087097606843912.087\,097\,606\,843\,912.087 097 606 843 91 7.410 347 403 810 2367.4103474038102367.410\,347\,403\,810\,2367.410 347 403 810 236 4.442 057 078 798 6364.4420570787986364.442\,057\,078\,798\,6364.442 057 078 798 636
1c Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}$ 17.264 941 603 848 42817.26494160384842817.264\,941\,603\,848\,42817.264 941 603 848 428 17.626 689 404 382 6417.6266894043826417.626\,689\,404\,382\,6417.626 689 404 382 64 2.432 730 621 280 0922.4327306212800922.432\,730\,621\,280\,0922.432 730 621 280 092 0.963 050 892 206 5430.9630508922065430.963\,050\,892\,206\,5430.963 050 892 206 543 3.304 831 562 414 722 43.30483156241472243.304\,831\,562\,414\,722\,43.304 831 562 414 722 4 3.187 231 883 395 354 33.18723188339535433.187\,231\,883\,395\,354\,33.187 231 883 395 354 3 3.366 519 174 041 2983.3665191740412983.366\,519\,174\,041\,2983.366 519 174 041 298 2.238 283 122 534 0172.2382831225340172.238\,283\,122\,534\,0172.238 283 122 534 017 7.209 562 943 593 5757.2095629435935757.209\,562\,943\,593\,5757.209 562 943 593 575 5.166 051 660 516 6055.1660516605166055.166\,051\,$6 051 6$ 605
2 Supervised S iNeuBe [51] - - - $15.1$ $0.954$
3a ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ 17.136 097 766 684 76317.13609776668476317.136\,097\,766\,684\,76317.136 097 766 684 763 17.561 826 082 102 69817.56182608210269817.561\,826\,082\,102\,69817.561 826 082 102 698 2.411 083 760 225 411 62.41108376022541162.411\,083\,760\,225\,411\,62.411 083 760 225 411 6 0.961 621 019 343 370 30.96162101934337030.961\,621\,019\,343\,370\,30.961 621 019 343 370 3 3.356 713 550 361 633 73.35671355036163373.356\,713\,550\,361\,633\,73.356 713 550 361 633 7 3.227 691 624 793 467 73.22769162479346773.227\,691\,624\,793\,467\,73.227 691 624 793 467 7 3.160 029 498 525 073 63.16002949852507363.160\,029\,498\,525\,073\,63.160 029 498 525 073 6 1.998 598 768 391 1651.9985987683911651.998\,598\,768\,391\,1651.998 598 768 391 165 6.840 679 865 521 1066.8406798655211066.840\,679\,865\,521\,1066.840 679 865 521 106 3.227 614 554 626 5593.2276145546265593.227\,614\,554\,626\,5593.227 614 554 626 559
3b ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}$ $\mathcal{L}_{X,q}^{\text{real}}$ 16.710 767 206 369 02416.71076720636902416.710\,767\,206\,369\,02416.710 767 206 369 024 17.254 252 569 589 13817.25425256958913817.254\,252\,569\,589\,13817.254 252 569 589 138 2.412 104 403 972 6262.4121044039726262.412\,104\,403\,972\,6262.412 104 403 972 626 0.960 142 528 725 129 80.96014252872512980.960\,142\,528\,725\,129\,80.960 142 528 725 129 8 3.322 187 319 009 8473.3221873190098473.322\,187\,319\,009\,8473.322 187 319 009 847 3.162 385 252 840 246 53.16238525284024653.162\,385\,252\,840\,246\,53.162 385 252 840 246 5 3.558 259 587 020 6493.5582595870206493.558\,259\,587\,020\,6493.558 259 587 020 649 2.083 410 155 241 712 32.08341015524171232.083\,410\,155\,241\,712\,32.083 410 155 241 712 3 6.854 688 083 675 7576.8546880836757576.854\,688\,083\,675\,7576.854 688 083 675 757 3.563 921 715 166 5193.5639217151665193.563\,921\,715\,166\,5193.563 921 715 166 519
3c ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}$ $\mathcal{L}_{X,q}^{\text{real}}$ 16.834 305 574 496 58616.83430557449658616.834\,305\,574\,496\,58616.834 305 574 496 586 17.251 441 973 969 0717.2514419739690717.251\,441\,973\,969\,0717.251 441 973 969 07 2.390 075 166 478 301 82.39007516647830182.390\,075\,166\,478\,301\,82.390 075 166 478 301 8 0.958 571 250 512 548 70.95857125051254870.958\,571\,250\,512\,548\,70.958 571 250 512 548 7 3.347 124 211 800 380 83.34712421180038083.347\,124\,211\,800\,380\,83.347 124 211 800 380 8 3.197 608 615 042 869 33.19760861504286933.197\,608\,615\,042\,869\,33.197 608 615 042 869 3 3.716 814 159 292 0353.7168141592920353.716\,814\,159\,292\,0353.716 814 159 292 035 2.072 347 800 435 119 32.07234780043511932.072\,347\,800\,435\,119\,32.072 347 800 435 119 3 7.639 148 300 336 1977.6391483003361977.639\,148\,300\,336\,1977.639 148 300 336 197 3.778 784 623 289 2713.7787846232892713.778\,784\,623\,289\,2713.778 784 623 289 271
4a ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $256$ 16.712 464 920 499 1116.7124649204991116.712\,464\,920\,499\,1116.712 464 920 499 11 17.167 910 347 776 17517.16791034777617517.167\,910\,347\,776\,17517.167 910 347 776 175 2.293 470 420 349 7282.2934704203497282.293\,470\,420\,349\,7282.293 470 420 349 728 0.957 785 015 609 969 50.95778501560996950.957\,785\,015\,609\,969\,50.957 785 015 609 969 5 3.332 327 093 435 673.332327093435673.332\,327\,093\,435\,673.332 327 093 435 67 3.190 962 144 995 6343.1909621449956343.190\,962\,144\,995\,6343.190 962 144 995 634 3.211 651 917 404 133.211651917404133.211\,651\,917\,404\,133.211 651 917 404 13 1.980 161 510 380 176 41.98016151038017641.980\,161\,510\,380\,176\,41.980 161 510 380 176 4 7.064 811 355 995 5187.0648113559955187.064\,811\,355\,995\,5187.064 811 355 995 518 3.409 780 933 252 373.409780933252373.409\,780\,933\,252\,373.409 780 933 252 37
4b ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $128$ 16.523 509 743 087 2216.5235097430872216.523\,509\,743\,087\,2216.523 509 743 087 22 17.081 067 725 176 98617.08106772517698617.081\,067\,725\,176\,98617.081 067 725 176 986 2.371 422 376 506 256 32.37142237650625632.371\,422\,376\,506\,256\,32.371 422 376 506 256 3 0.959 743 330 086 607 50.95974333008660750.959\,743\,330\,086\,607\,50.959 743 330 086 607 5 3.354 864 545 712 0263.3548645457120263.354\,864\,545\,712\,0263.354 864 545 712 026 3.224 684 022 093 959 43.22468402209395943.224\,684\,022\,093\,959\,43.224 684 022 093 959 4 3.407 079 646 017 699 43.40707964601769943.407\,079\,646\,017\,699\,43.407 079 646 017 699 4 1.876 912 865 518 641.876912865518641.876\,912\,865\,518\,641.876 912 865 518 64 7.083 488 980 201 7197.0834889802017197.083\,488\,980\,201\,7197.083 488 980 201 719 3.199 588 957 914 8963.1995889579148963.199\,588\,957\,914\,8963.199 588 957 914 896
4c ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $64$ 17.016 686 683 125 56717.01668668312556717.016\,686\,683\,125\,56717.016 686 683 125 567 17.537 237 975 786 34317.53723797578634317.537\,237\,975\,786\,34317.537 237 975 786 343 2.402 777 045 513 644 32.40277704551364432.402\,777\,045\,513\,644\,32.402 777 045 513 644 3 0.961 299 180 774 926 30.96129918077492630.961\,299\,180\,774\,926\,30.961 299 180 774 926 3 3.380 592 781 113 3393.3805927811133393.380\,592\,781\,113\,3393.380 592 781 113 339 3.272 828 277 426 903 73.27282827742690373.272\,828\,277\,426\,903\,73.272 828 277 426 903 7 3.289 085 545 722 7143.2890855457227143.289\,085\,545\,722\,7143.289 085 545 722 714 1.869 537 962 314 244 51.86953796231424451.869\,537\,962\,314\,244\,51.869 537 962 314 244 5 6.966 753 828 912 9626.9667538289129626.966\,753\,828\,912\,9626.966 753 828 912 962 3.134 195 898 921 0153.1341958989210153.134\,195\,898\,921\,0153.134 195 898 921 015
4d ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $32$ 16.478 895 149 899 2516.4788951498992516.478\,895\,149\,899\,2516.478 895 149 899 25 17.141 230 481 170 06617.14123048117006617.141\,230\,481\,170\,06617.141 230 481 170 066 2.360 100 678 873 7852.3601006788737852.360\,100\,678\,873\,7852.360 100 678 873 785 0.961 100 110 551 726 20.96110011055172620.961\,100\,110\,551\,726\,20.961 100 110 551 726 2 3.345 090 992 793 603 23.34509099279360323.345\,090\,992\,793\,603\,23.345 090 992 793 603 2 3.224 999 633 733 2563.2249996337332563.224\,999\,633\,733\,2563.224 999 633 733 256 3.384 955 752 212 389 53.38495575221238953.384\,955\,752\,212\,389\,53.384 955 752 212 389 5 1.987 536 413 584 571 61.98753641358457161.987\,536\,413\,584\,571\,61.987 536 413 584 571 6 6.957 415 016 809 8626.9574150168098626.957\,415\,016\,809\,8626.957 415 016 809 862 3.545 237 984 025 413.545237984025413.545\,237\,984\,025\,413.545 237 984 025 41
4e ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $16$ 16.078 147 192 231 63316.07814719223163316.078\,147\,192\,231\,63316.078 147 192 231 633 16.928 411 923 202 48216.92841192320248216.928\,411\,923\,202\,48216.928 411 923 202 482 2.364 959 328 012 0332.3649593280120332.364\,959\,328\,012\,0332.364 959 328 012 033 0.959 443 197 977 285 30.95944319797728530.959\,443\,197\,977\,285\,30.959 443 197 977 285 3 3.361 895 135 599 697 33.36189513559969733.361\,895\,135\,599\,697\,33.361 895 135 599 697 3 3.244 502 963 142 3823.2445029631423823.244\,502\,963\,142\,3823.244 502 963 142 382 3.436 578 171 091 4463.4365781710914463.436\,578\,171\,091\,4463.436 578 171 091 446 1.965 411 703 971 385 21.96541170397138521.965\,411\,703\,971\,385\,21.965 411 703 971 385 2 7.228 240 567 799 7767.2282405677997767.228\,240\,567\,799\,7767.228 240 567 799 776 3.040 777 243 215 473.040777243215473.040\,777\,243\,215\,473.040 777 243 215 47
TABLE IV: ctPuLSE vs. Purely-Supervised Models on CHiME-4 Far-Field Mixtures (Six-Channel Input)
[…]
0 Mixture - - - - - 7.506 217 692 752 9057.5062176927529057.506\,217\,692\,752\,9057.506 217 692 752 905 7.539 180 193 284 8817.5391801932848817.539\,180\,193\,284\,8817.539 180 193 284 881 1.273 397 988 803 4591.2733979888034591.273\,397\,988\,803\,4591.273 397 988 803 459 0.870 243 137 103 452 40.87024313710345240.870\,243\,137\,103\,452\,40.870 243 137 103 452 4 1.520 743 794 895 107 71.52074379489510771.520\,743\,794\,895\,107\,71.520 743 794 895 107 7 1.392 663 582 045 257 81.39266358204525781.392\,663\,582\,045\,257\,81.392 663 582 045 257 8 5.932 890 8555.9328908555.932\,890\,8555.932 890 855 4.070 946 5684.0709465684.070\,946\,5684.070 946 568 8.288 195 7418.2881957418.288\,195\,7418.288 195 741 4.470 082 6754.4700826754.470\,082\,6754.470 082 675
1a Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ 22.913 916 111 895 5322.9139161118955322.913\,916\,111\,895\,5322.913 916 111 895 53 23.322 007 247 411 8523.3220072474118523.322\,007\,247\,411\,8523.322 007 247 411 85 3.344 010 585 185 2533.3440105851852533.344\,010\,585\,185\,2533.344 010 585 185 253 0.987 202 710 663 9770.9872027106639770.987\,202\,710\,663\,9770.987 202 710 663 977 2.120 078 131 527 103 32.12007813152710332.120\,078\,131\,527\,103\,32.120 078 131 527 103 3 1.776 820 183 864 4611.7768201838644611.776\,820\,183\,864\,4611.776 820 183 864 461 0.862 831 858 407 079 60.86283185840707960.862\,831\,858\,407\,079\,60.862 831 858 407 079 6 23.164 570 965 006 08423.16457096500608423.164\,570\,965\,006\,08423.164 570 965 006 084 1.340 119 536 794 919 71.34011953679491971.340\,119\,536\,794\,919\,71.340 119 536 794 919 7 46.713 998 785 557 47646.71399878555747646.713\,998\,785\,557\,47646.713 998 785 557 476
1b Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}$ 22.955 472 425 077 922.955472425077922.955\,472\,425\,077\,922.955 472 425 077 9 23.339 479 570 583 49623.33947957058349623.339\,479\,570\,583\,49623.339 479 570 583 496 3.289 384 816 090 2663.2893848160902663.289\,384\,816\,090\,2663.289 384 816 090 266 0.987 573 048 572 883 20.98757304857288320.987\,573\,048\,572\,883\,20.987 573 048 572 883 2 2.063 506 082 606 499 32.06350608260649932.063\,506\,082\,606\,499\,32.063 506 082 606 499 3 1.631 255 842 475 493 51.63125584247549351.631\,255\,842\,475\,493\,51.631 255 842 475 493 5 0.855 457 227 138 6430.8554572271386430.855\,457\,227\,138\,6430.855 457 227 138 643 40.270 658 947 601 3140.2706589476013140.270\,658\,947\,601\,3140.270 658 947 601 31 1.316 772 506 537 168 51.31677250653716851.316\,772\,506\,537\,168\,51.316 772 506 537 168 5 64.711 102 807 230 664.711102807230664.711\,102\,807\,230\,664.711 102 807 230 6
1c Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}$ 22.809 102 938 030 70522.80910293803070522.809\,102\,938\,030\,70522.809 102 938 030 705 23.203 917 606 250 7823.2039176062507823.203\,917\,606\,250\,7823.203 917 606 250 78 3.215 278 488 668 528 43.21527848866852843.215\,278\,488\,668\,528\,43.215 278 488 668 528 4 0.987 270 892 267 386 20.98727089226738620.987\,270\,892\,267\,386\,20.987 270 892 267 386 2 2.043 228 114 465 531 82.04322811446553182.043\,228\,114\,465\,531\,82.043 228 114 465 531 8 1.627 247 462 232 174 21.62724746223217421.627\,247\,462\,232\,174\,21.627 247 462 232 174 2 0.825 958 702 064 896 70.82595870206489670.825\,958\,702\,064\,896\,70.825 958 702 064 896 7 53.541 797 263 910 9153.5417972639109153.541\,797\,263\,910\,9153.541 797 263 910 91 1.298 094 882 330 967 51.29809488233096751.298\,094\,882\,330\,967\,51.298 094 882 330 967 5 74.216 451 025 269 7574.2164510252697574.216\,451\,025\,269\,7574.216 451 025 269 75
2a Supervised S iNeuBe [51] - - - $22.0$ $22.4$ $0.986$
2b Supervised S SpatialNet [52] - - - $22.1$ $22.3$ $2.88$ $0.983$
2c Supervised S USES [12] - - - - $20.6$ $3.16$ $0.983$ $4.2$ $78.1$
2d Supervised S USES2 [13] - - - - $18.8$ $2.94$ $0.979$ $2.96$ $4.6$ $12.1$
3a ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ 22.551 012 031 959 76522.55101203195976522.551\,012\,031\,959\,76522.551 012 031 959 765 22.848 970 369 805 5322.8489703698055322.848\,970\,369\,805\,5322.848 970 369 805 53 3.114 229 646 144 4913.1142296461444913.114\,229\,646\,144\,4913.114 229 646 144 491 0.984 923 755 300 023 90.98492375530002390.984\,923\,755\,300\,023\,90.984 923 755 300 023 9 3.384 385 940 797 4793.3843859407974793.384\,385\,940\,797\,4793.384 385 940 797 479 3.278 956 317 309 036 43.27895631730903643.278\,956\,317\,309\,036\,43.278 956 317 309 036 4 0.818 584 070 796 460 20.81858407079646020.818\,584\,070\,796\,460\,20.818 584 070 796 460 2 1.283 233 157 564 8071.2832331575648071.283\,233\,157\,564\,8071.283 233 157 564 807 1.391 483 003 361 972 31.39148300336197231.391\,483\,003\,361\,972\,31.391 483 003 361 972 3 1.653 510 205 988 1361.6535102059881361.653\,510\,205\,988\,1361.653 510 205 988 136
3b ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}$ $\mathcal{L}_{X,q}^{\text{real}}$ 22.805 953 327 453 49822.80595332745349822.805\,953\,327\,453\,49822.805 953 327 453 498 23.274 481 308 482 4523.2744813084824523.274\,481\,308\,482\,4523.274 481 308 482 45 3.262 422 973 278 7973.2624229732787973.262\,422\,973\,278\,7973.262 422 973 278 797 0.986 982 073 716 504 10.98698207371650410.986\,982\,073\,716\,504\,10.986 982 073 716 504 1 3.389 140 097 976 963.389140097976963.389\,140\,097\,976\,963.389 140 097 976 96 3.281 604 502 545 969 53.28160450254596953.281\,604\,502\,545\,969\,53.281 604 502 545 969 5 0.870 206 489 675 516 20.87020648967551620.870\,206\,489\,675\,516\,20.870 206 489 675 516 2 1.272 170 802 758 213 81.27217080275821381.272\,170\,802\,758\,213\,81.272 170 802 758 213 8 1.349 458 348 898 020 11.34945834889802011.349\,458\,348\,898\,020\,11.349 458 348 898 020 1 1.732 916 063 337 848 51.73291606333784851.732\,916\,063\,337\,848\,51.732 916 063 337 848 5
3c ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}$ $\mathcal{L}_{X,q}^{\text{real}}$ 22.795 234 206 047 92622.79523420604792622.795\,234\,206\,047\,92622.795 234 206 047 926 23.116 705 751 770 43623.11670575177043623.116\,705\,751\,770\,43623.116 705 751 770 436 3.289 739 032 857 345 73.28973903285734573.289\,739\,032\,857\,345\,73.289 739 032 857 345 7 0.986 985 045 099 122 80.98698504509912280.986\,985\,045\,099\,122\,80.986 985 045 099 122 8 3.394 377 605 833 2813.3943776058332813.394\,377\,605\,833\,2813.394 377 605 833 281 3.302 837 401 825 815 53.30283740182581553.302\,837\,401\,825\,815\,53.302 837 401 825 815 5 0.844 395 280 235 988 30.84439528023598830.844\,395\,280\,235\,988\,30.844 395 280 235 988 3 1.286 920 609 167 004 81.28692060916700481.286\,920\,609\,167\,004\,81.286 920 609 167 004 8 1.298 094 882 330 967 51.29809488233096751.298\,094\,882\,330\,967\,51.298 094 882 330 967 5 1.564 762 483 067 868 81.56476248306786881.564\,762\,483\,067\,868\,81.564 762 483 067 868 8
4a ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $256$ 22.646 991 805 596 78422.64699180559678422.646\,991\,805\,596\,78422.646 991 805 596 784 22.958 123 369 688 53322.95812336968853322.958\,123\,369\,688\,53322.958 123 369 688 533 3.112 808 776 714 5853.1128087767145853.112\,808\,776\,714\,5853.112 808 776 714 585 0.985 584 073 681 268 40.98558407368126840.985\,584\,073\,681\,268\,40.985 584 073 681 268 4 3.367 654 219 370 248 63.36765421937024863.367\,654\,219\,370\,248\,63.367 654 219 370 248 6 3.249 435 543 997 756 23.24943554399775623.249\,435\,543\,997\,756\,23.249 435 543 997 756 2 0.855 457 227 138 6430.8554572271386430.855\,457\,227\,138\,6430.855 457 227 138 643 1.305 357 867 177 993 31.30535786717799331.305\,357\,867\,177\,993\,31.305 357 867 177 993 3 1.340 119 536 794 919 71.34011953679491971.340\,119\,536\,794\,919\,71.340 119 536 794 919 7 1.770 283 525 620 066 51.77028352562006651.770\,283\,525\,620\,066\,51.770 283 525 620 066 5
4b ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $128$ 22.877 807 343 547 99622.87780734354799622.877\,807\,343\,547\,99622.877 807 343 547 996 23.222 930 376 692 1323.2229303766921323.222\,930\,376\,692\,1323.222 930 376 692 13 3.219 571 926 647 966 53.21957192664796653.219\,571\,926\,647\,966\,53.219 571 926 647 966 5 0.986 151 257 697 490 80.98615125769749080.986\,151\,257\,697\,490\,80.986 151 257 697 490 8 3.340 576 934 824 8893.3405769348248893.340\,576\,934\,824\,8893.340 576 934 824 889 3.251 119 989 958 0863.2511199899580863.251\,119\,989\,958\,0863.251 119 989 958 086 0.837 020 648 967 551 70.83702064896755170.837\,020\,648\,967\,551\,70.837 020 648 967 551 7 1.323 795 125 188 9821.3237951251889821.323\,795\,125\,188\,9821.323 795 125 188 982 1.335 450 130 743 369 41.33545013074336941.335\,450\,130\,743\,369\,41.335 450 130 743 369 4 1.700 219 533 840 9081.7002195338409081.700\,219\,533\,840\,9081.700 219 533 840 908
4c ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $64$ 22.804 458 647 966 38522.80445864796638522.804\,458\,647\,966\,38522.804 458 647 966 385 23.199 798 023 054 06523.19979802305406523.199\,798\,023\,054\,06523.199 798 023 054 065 3.199 864 766 453 2523.1998647664532523.199\,864\,766\,453\,2523.199 864 766 453 252 0.986 137 138 983 365 20.98613713898336520.986\,137\,138\,983\,365\,20.986 137 138 983 365 2 3.385 378 211 815 196 33.38537821181519633.385\,378\,211\,815\,196\,33.385 378 211 815 196 3 3.289 810 700 044 8533.2898107000448533.289\,810\,700\,044\,8533.289 810 700 044 853 0.859 144 542 772 861 40.85914454277286140.859\,144\,542\,772\,861\,40.859 144 542 772 861 4 1.253 733 544 747 225 21.25373354474722521.253\,733\,544\,747\,225\,21.253 733 544 747 225 2 1.368 135 973 104 221 21.36813597310422121.368\,135\,973\,104\,221\,21.368 135 973 104 221 2 1.700 219 533 840 9081.7002195338409081.700\,219\,533\,840\,9081.700 219 533 840 908
4d ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $32$ 22.713 268 510 319 9722.7132685103199722.713\,268\,510\,319\,9722.713 268 510 319 97 23.082 258 429 569 07723.08225842956907723.082\,258\,429\,569\,07723.082 258 429 569 077 3.234 548 879 572 839 73.23454887957283973.234\,548\,879\,572\,839\,73.234 548 879 572 839 7 0.986 062 905 656 321 20.98606290565632120.986\,062\,905\,656\,321\,20.986 062 905 656 321 2 3.391 553 352 167 8393.3915533521678393.391\,553\,352\,167\,8393.391 553 352 167 839 3.302 316 856 079 836 63.30231685607983663.302\,316\,856\,079\,836\,63.302 316 856 079 836 6 0.870 206 489 675 516 20.87020648967551620.870\,206\,489\,675\,516\,20.870 206 489 675 516 2 1.279 545 705 962 609 21.27954570596260921.279\,545\,705\,962\,609\,21.279 545 705 962 609 2 1.307 433 694 434 0681.3074336944340681.307\,433\,694\,434\,0681.307 433 694 434 068 1.672 193 937 129 244 61.67219393712924461.672\,193\,937\,129\,244\,61.672 193 937 129 244 6
4e ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $16$ 22.581 081 933 595 91822.58108193359591822.581\,081\,933\,595\,91822.581 081 933 595 918 22.957 867 337 549 6922.9578673375496922.957\,867\,337\,549\,6922.957 867 337 549 69 3.245 428 087 765 523.245428087765523.245\,428\,087\,765\,523.245 428 087 765 52 0.985 162 009 465 4390.9851620094654390.985\,162\,009\,465\,4390.985 162 009 465 439 3.379 546 646 573 118 53.37954664657311853.379\,546\,646\,573\,118\,53.379 546 646 573 118 5 3.273 793 347 574 7313.2737933475747313.273\,793\,347\,574\,7313.273 793 347 574 731 0.888 643 067 846 607 70.88864306784660770.888\,643\,067\,846\,607\,70.888 643 067 846 607 7 1.261 108 447 951 620 61.26110844795162061.261\,108\,447\,951\,620\,61.261 108 447 951 620 6 1.419 499 439 671 273 81.41949943967127381.419\,499\,439\,671\,273\,81.419 499 439 671 273 8 1.798 309 122 331 729 51.79830912233172951.798\,309\,122\,331\,729\,51.798 309 122 331 729 5
VII-B Key Results of Far-Field Speech Enhancement
Table III and IV respectively present the results of monaural and six-channel ctPuLSEnet on CHiME-4.
Let us first provide the hyper-parameter configurations of ctPuLSEnet. The frequency-domain linear filter in (6) is tuned to $1$-tap (i.e., in the text below (6), $I=1$ and $J=0$). In (8), the time-domain filter tap $K$ is tuned based on the set of $\{256,128,64,32,16\}$. In this case, the linear filter length $K+1+K\in\{513,257,129,65,33\}$ is comparable to or shorter than the STFT window length, which is $512$ samples long in this paper (see the STFT configurations in Section VI-D). The weighting term $\alpha$ in (14) is tuned to $5$. In both tables, when the filter tap for time-domain linear filtering (i.e., the “$K$” column) is denoted as “-”, it means that frequency-domain linear filtering in (6) is used. Otherwise, time-domain linear filtering in (9) is used.
TABLE V: Comparison of ctPuLSE with Other Approaches on CHiME-4 Far-Field Mixtures
[…]
Cross reference Joint training? #input mics Val. Test Val. Test
Row Systems Frontend REAL REAL SIMU REAL SIMU REAL
0 - Mixture - - 1 - 1.520 743 794 895 107 71.52074379489510771.520\,743\,794\,895\,107\,71.520 743 794 895 107 7 1.392 663 582 045 257 81.39266358204525781.392\,663\,582\,045\,257\,81.392 663 582 045 257 8 5.932 890 8555.9328908555.932\,890\,8555.932 890 855 4.070 946 5684.0709465684.070\,946\,5684.070 946 568 8.288 195 7418.2881957418.288\,195\,7418.288 195 741 4.470 082 6754.4700826754.470\,082\,6754.470 082 675
1a - IRIS [44] Conv-TasNet 1 - - - $5.96$ $4.37$ $13.52$ $12.11$
1b - IRIS [44] Conv-TasNet 1 - - - $3.16$ $2.03$ $6.12$ $3.92$
2 - SuperM2M [3] TF-GridNet 1 - 3.163 384 205 610 3933.1633842056103933.163\,384\,205\,610\,3933.163 384 205 610 393 3.033 267 061 570 5773.0332670615705773.033\,267\,061\,570\,5773.033 267 061 570 577 3.392 330 3833.3923303833.392\,330\,3833.392 330 383 1.836 350 8971.8363508971.836\,350\,8971.836 350 897 6.569 854 3146.5698543146.569\,854\,3146.569 854 314 3.036 106 313.036106313.036\,106\,313.036 106 31
[…]
4a - MultiIRIS [46] Neural WPD 2 - - - $2.28$ $2.06$ $2.30$ $3.63$
4b - MultiIRIS [46] Neural WPD 2 - - - $2.04$ $1.66$ $2.04$ $2.65$
5 - SuperM2M [3] TF-GridNet 2 - 3.006 918 163 792 940 33.00691816379294033.006\,918\,163\,792\,940\,33.006 918 163 792 940 3 2.903 423 564 541 1312.9034235645411312.903\,423\,564\,541\,1312.903 423 564 541 131 1.497 050 1471.4970501471.497\,050\,1471.497 050 147 1.401 231 6081.4012316081.401\,231\,6081.401 231 608 2.082 555 0982.0825550982.082\,555\,0982.082 555 098 1.943 108 0381.9431080381.943\,108\,0381.943 108 038
[…]
7a - MultiIRIS [46] Neural WPD 6 - - - $1.19$ $1.32$ $1.29$ $1.85$
7b - MultiIRIS [46] Neural WPD 6 - - - $1.22$ $1.33$ $1.24$ $1.77$
8 - SuperM2M [3] TF-GridNet 6 - 2.839 865 817 670 8372.8398658176708372.839\,865\,817\,670\,8372.839 865 817 670 837 2.749 583 063 702 629 72.74958306370262972.749\,583\,063\,702\,629\,72.749 583 063 702 629 7 0.829 646 0170.8296460170.829\,646\,0170.829 646 017 1.257 420 9961.2574209961.257\,420\,9961.257 420 996 1.368 135 9731.3681359731.368\,135\,9731.368 135 973 1.606 800 8781.6068008781.606\,800\,8781.606 800 878
[…]
Note #$\#1$: The best scores are highlighted in bold in each of the $1$-, $2$- and $6$-channel tasks separately.
Note #$\#2$: SuperM2M and ctPuLSE use exactly the same TF-GridNet architecture.
Note #$\#3$: The “cross reference” entry denotes that the other configurations are the same as the ones in the referred row in another table.
The key results of this paper are presented in row 00, $1$a and $3$a of Table III and IV. Comparing row $1$a with 00, we observe that, on the simulated test data, purely-supervised learning based TF-GridNet trained on simulated data obtains strong enhancement results [^4] By “enhancement results”, we mean SI-SDR, SDR, WB-PESQ and STOI scores. (e.g., $17.3$ vs. $7.5$ dB SI-SDR in Table III, and $22.9$ vs. $7.5$ dB SI-SDR in Table IV), and strong ASR performance (e.g., $7.63$% vs. $8.29$% WER in Table III, and $1.34$% vs. $8.29$% WER in Table IV). In addition, the enhancement results on the simulated test data obtained by TF-GridNet in row $1$a are better than strong existing models such as iNeuBe [51, 53], SpatialNet [52], USES [12] and USES2 [13]. However, the performance on the real test data is limited. For example, the ASR performance on the real test data is degraded compared to just using unprocessed mixtures for ASR decoding (e.g., $4.47$% vs. $5.24$% WER in Table III). This degradation is much more severe in multi-channel cases (e.g., $4.47$% vs. $46.71$% WER in Table IV), possibly because simulated inter-microphone characteristics are more likely to mismatch those in real-recorded multi-channel data, compared with monaural cases where only one microphone needs to be simulated. These problems are widely-observed in earlier robust ASR studies [11, 23], largely because (a) on real data, enhancement models trained on simulated data usually introduce speech distortion detrimental to ASR; and (b) simulated training data is often mismatched with real-recorded test data. In row $3$a, our proposed ctPuLSEnet, trained on simulated and real data combined (denoted as S+R in the “Training data” column), obtains clearly better ASR performance on the real test set over row $1$a and 00 (e.g., $3.23$% vs. $5.24$% and $4.47$% in Table III, and $1.65$% vs. $46.71$% and $4.47$% in Table IV). These results indicate that ctPuLSE is an effective mechanism for learning from real-recorded data, and can yield enhancement models with better generalizability to real data than purely-supervised approaches which train enhancement models only on simulated data.
VII-C Ablation Results of Far-Field Speech Enhancement
Next, we present several ablation results of ctPuLSEnet.
Comparing row $3$a with $3$b and $3$c in Table III, we observe that configuring ctPuLSEnet to estimate noise besides target speech and at the same time including both the loss on the noise estimate and the mixture-constraint loss (i.e., row $3$a) lead to better enhancement and ASR performance. Comparing $3$a with $1$a, $3$b with $1$b, and $3$c with $1$c, we observe that ctPuLSE obtains better ASR performance on the real test data even if various loss functions are used for the simulated data.
In $4$a-$4$e of Table III, we switch from frequency-domain linear filtering to time-domain linear filtering, and experiment with various filter lengths by tuning $K$ (defined in the text below (8)) based on the set of $\{256,128,64,32,16\}$. We observe that setting $K$ to $64$ (in row $4$c) produces the best ASR performance on the real validation set among various options, and the ASR performance is also better than row $3$a on the real validation set (i.e., $1.87$% vs. $2.00$% WER in Table III). Similar trend is also observed in the six-channel-input case in Table IV. We therefore choose the setup in row $4$c for the rest of experiments in this paper.
VII-D Comparison with Other Approaches
In Table VII-B, we compare the performance of ctPuLSEnet with IRIS [44], multi-channel IRIS (MultiIRIS) [46] and SuperM2M [3] on the $1$-, $2$-, and $6$-channel tasks of CHiME-4.
IRIS and MultiIRIS were the state-of-the-art systems on CHiME-4 before SuperM2M. Comparing row $1$a with $1$b, $4$a with $4$b, and $7$a with $7$b, we observe that IRIS and MultiIRIS need joint frontend-backend training to achieve strong ASR performance, especially in the $1$- and $2$-channel cases.
SuperM2M [3], even without joint frontend-backend training, achieves better ASR performance on the real data than IRIS and MultiIRIS. However, the enhancement results (measured by DNSMOS OVRL) of SuperM2M on the real data are not strong. Upon listening to its processed signals [^5] See a sound demo at https://zqwang7.github.io/demos/ctPuLSE_demo/index.html, which provides a comparison between SuperM2M and ctPuLSE., we observe that it cannot suppress noises sufficiently and tends to maintain some noise signals in its estimate of target speech, especially in multi-channel cases. This is reflected by the DNSMOS OVRL scores in row $2$, $5$ and $8$, and it is quite counter-intuitive that the score decreases when the number of input microphones increases (e.g., from $3.03$ in the monaural case down to $2.90$ in the $2$-channel case and down to $2.75$ in the $6$-channel case on the real test data). We think that the mediocre enhancement score is likely because, in SuperM2M [3], the loss function is defined on observed mixtures (i.e., between each observed mixture and reconstructed mixture, which is obtained by summing up linearly-filtered source estimates) rather than on individual source estimates. In this case, the DNN would only have a weak supervision regarding what the target sources are. That is, the cues leveraged for training enhancement models are (a) there should be two sources; and (b) their estimates, after linear filtering, should add up to each mixture. Such a supervision could be too weak for the DNN to achieve good enhancement.
In comparison, in ctPuLSE, the loss function includes a loss term on the pseudo-target speech provided by CTSEnet (i.e., $\mathcal{L}_{X}^{\text{real}}$ in (5)). As long as the pseudo-target is reasonably accurate, ctPuLSE is expected to more accurately suppress non-target signals. Comparing row $3$ with $2$, $6$ with $5$, and $9$ with $8$, we observe that ctPuLSE indeed obtains dramatically better DNSMOS OVRL scores than SuperM2M, and the enhanced target speech sounds much cleaner (please see the sound demo). Although the ASR performance on the real test data is worse, it is only slightly worse and is still very strong.
VII-E Miscellaneous Results
Table VII-E reports the results of applying speaker reinforcement (see the details in Section VI-H), where the SNR factor $\gamma$ is tuned to $10$ dB. We observe that the performance gap between SuperM2M and ctPuLSE observed in Table VII-B is reduced, especially on the $2$- and $6$-channel tasks. Our best system in row $6$b obtains ASR results competitive to using close-talk mixtures for ASR decoding, suggesting the effectiveness of ctPuLSE and the overall robust ASR system.
TABLE VI: Effects of Speaker Reinforcement
Spk. reinf. $\gamma$ (dB) WER (%$\downarrow$
#input mics Val. Test
Row Systems SIMU REAL SIMU REAL
0 Mixture 1 - 5.932 890 8555.9328908555.932\,890\,8555.932 890 855 4.070 946 5684.0709465684.070\,946\,5684.070 946 568 8.288 195 7418.2881957418.288\,195\,7418.288 195 741 4.470 082 6754.4700826754.470\,082\,6754.470 082 675
1a SuperM2M [3] 1 - 3.392 330 3833.3923303833.392\,330\,3833.392 330 383 1.836 350 8971.8363508971.836\,350\,8971.836 350 897 6.569 854 3146.5698543146.569\,854\,3146.569 854 314 3.036 106 313.036106313.036\,106\,313.036 106 31
1b SuperM2M [3] 1 $10$ 2.400 442 4772.4004424772.400\,442\,4772.400 442 477 1.640 915 9621.6409159621.640\,915\,9621.640 915 962 4.538 662 6824.5386626824.538\,662\,6824.538 662 682 2.400 859 4512.4008594512.400\,859\,4512.400 859 451
[…]
2b ctPuLSE 1 $10$ 2.293 510 324 483 7762.2935103244837762.293\,510\,324\,483\,7762.293 510 324 483 776 1.666 728 124 193 371.666728124193371.666\,728\,124\,193\,371.666 728 124 193 37 4.758 124 766 529 6984.7581247665296984.758\,124\,766\,529\,6984.758 124 766 529 698 2.550 329 300 761 3622.5503293007613622.550\,329\,300\,761\,3622.550 329 300 761 362
3a SuperM2M [3] 2 - 1.497 050 1471.4970501471.497\,050\,1471.497 050 147 1.401 231 6081.4012316081.401\,231\,6081.401 231 608 2.082 555 0982.0825550982.082\,555\,0982.082 555 098 1.943 108 0381.9431080381.943\,108\,0381.943 108 038
3b SuperM2M [3] 2 $10$ 1.283 185 841.283185841.283\,185\,841.283 185 84 1.331 170 0281.3311700281.331\,170\,0281.331 170 028 1.881 770 6381.8817706381.881\,770\,6381.881 770 638 1.835 676 5841.8356765841.835\,676\,5841.835 676 584
[…]
4b ctPuLSE 2 $10$ 1.445 427 728 613 569 31.44542772861356931.445\,427\,728\,613\,569\,31.445 427 728 613 569 3 1.356 982 189 608 761 31.35698218960876131.356\,982\,189\,608\,761\,31.356 982 189 608 761 3 2.003 175 196 115 0542.0031751961150542.003\,175\,196\,115\,0542.003 175 196 115 054 1.845 018 450 184 5021.8450184501845021.845\,018\,450\,184\,5021.845 018 450 184 502
5a SuperM2M [3] 6 - 0.829 646 0170.8296460170.829\,646\,0170.829 646 017 1.257 420 9961.2574209961.257\,420\,9961.257 420 996 1.368 135 9731.3681359731.368\,135\,9731.368 135 973 1.606 800 8781.6068008781.606\,800\,8781.606 800 878
5b SuperM2M [3] 6 $10$ 0.833 333 3330.8333333330.833\,333\,3330.833 333 333 1.231 608 8351.2316088351.231\,608\,8351.231 608 835 1.372 805 3791.3728053791.372\,805\,3791.372 805 379 1.578 775 2811.5787752811.578\,775\,2811.578 775 281
[…]
6b ctPuLSE 6 $10$ 0.855 457 227 138 6430.8554572271386430.855\,457\,227\,138\,6430.855 457 227 138 643 1.224 233 931 929 643 41.22423393192964341.224\,233\,931\,929\,643\,41.224 233 931 929 643 4 1.372 805 379 155 771 41.37280537915577141.372\,805\,379\,155\,771\,41.372 805 379 155 771 4 1.564 762 483 067 868 81.56476248306786881.564\,762\,483\,067\,868\,81.564 762 483 067 868 8
7 Close-talk Mixture - - - 1.143 109 9961.1431099961.143\,109\,9961.143 109 996 - 1.490 027 5581.4900275581.490\,027\,5581.490 027 558
Note: Oracle results (i.e., directly using close-talk mixtures for ASR decoding)
       are marked in grey.
VIII Conclusion
We have investigated close-talk speech enhancement, and have proposed a novel approach, ctPuLSE, for far-field speech enhancement, where estimated close-talk speech produced by close-talk speech enhancement is utilized as pseudo-labels for training supervised enhancement models directly on real far-field mixtures to realize better generalizability to real data. Evaluation results on the challenging CHiME-4 dataset show the effectiveness and potential of the proposed algorithms.
Although simple and straightforward, the proposed approach of exploiting close-talk mixtures for far-field speech enhancement, we think, could encourage a new stream of research towards realizing neural speech enhancement models that can generalize better to real data, as it suggests a promising way that can derive, for real mixtures, pseudo-labels which can enable the training of speech enhancement models directly on real mixtures. Thanks to the innate high input SNR of close-talk mixtures, the derived pseudo-labels are often reliable and high-quality. This paper, based on the challenging CHiME-4 dataset, has shown that the derived pseudo-labels can be utilized to build far-field speech enhancement models with better generalizability to real data. Looking forward, we expect the derived pseudo-labels to be also useful in many other applications beyond far-field speech enhancement.